OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi /10

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

coterminal means repeating at full \(360^{\circ}\) so \(\dfrac{π}{10}=18^{\circ}\) thus coterminals of \(\dfrac{π}{10}=18^{\circ}\) are \(18^{\circ}±k ~360^{\circ}\) substitute \( k=1\) coterminals are \(18+360=540^{\circ}~~ and~~ 18−360=-342^{\circ}\)